OpenStudy (anonymous):

f(x) = x2 + 3; g(x) = sqrt(x - 2) Find f(g(x)).

OpenStudy (jhannybean):

\[\large f(x) = x^2 +3\]\[\large g(x) = \sqrt{x-2}\] so replace all the x's in f(x) with the function of g(x). \[\large f(g(x)) = (\sqrt{x-2})^2 +3 \] When you square a square root, the square root and the square cancel eachother out.

OpenStudy (jhannybean):

Can you finish solving this now? :)

OpenStudy (jhannybean):

They don't necessarily cancel each other out, but when you take the product of two square roots, it becomes an integer.

OpenStudy (anonymous):

i might be helpful to take an intermediate step maybe not \[f(g(x))=f(\sqrt{x-2})\] may help

OpenStudy (anonymous):

replace the general \(g(x)\) by the specific one you have in this case then replace the \(x\) in \(f(x)=x^2+3\) by \(\sqrt{x-2}\)

OpenStudy (anonymous):

also helpful not to think of \(x\) when you do this \[f(\diamondsuit)=\diamondsuit^2 +3\] replace \(\diamondsuit\) by \(\sqrt{x-2}\)

OpenStudy (anonymous):

x+1 answer